# Inland-LOAM: Voxel-Based Structural Semantic Mapping for Inland Waterways

### v1

@@ Abstract
Accurate and up-to-date geospatial information is crucial for enhancing the safety and autonomy of Inland Waterway Transport (IWT). Inland Electronic Navigational Charts (IENC) often lack detailed real-time information, conventional LiDAR-based Simultaneous Localization and Mapping (SLAM) methods have difficulty handling distinctive inland waterway characteristics, such as dynamic shorelines and challenging reflective water surfaces. These factors lead to significant localization drift, particularly in the vertical direction, and produce dense point cloud maps without the semantic richness required for effective autonomous decision making.
@@
This paper introduces Inland-LOAM, a comprehensive LiDAR-based SLAM framework specifically for inland waterway environments. Firstly, we present an improved feature extraction method adapted to the unique geometric characteristics of waterways, combined with a joint optimization framework that incorporates the water surface as a global planar constraint to significantly mitigate vertical drift. Secondly, we propose an innovative pipeline to transform dense 3D point cloud outputs into structured 2D semantic maps by constructing semantic voxel grids and performing geometric analyses (roughness, planarity, and slope). This enables classification into structurally meaningful categories and supports real-time computation of critical navigational parameters such as vertical bridge clearances. Additionally, an automated map conversion module efficiently extracts shoreline boundaries via ray-casting and exports these features into a lightweight, IENC-compatible format.
Extensive evaluations were conducted using a diverse, real-world dataset featuring urban canals, vegetated banks, and bridge crossings. Comparative results demonstrate that Inland-LOAM achieves superior localization accuracy over state-of-the-art LiDAR SLAM methods. Generated semantic maps and extracted shoreline boundaries align with real-world conditions, as verified through comparisons with reference maps, thus providing reliable and up-to-date geospatial information to enhance navigational situational awareness. Both the dataset and the algorithm are publicly available to support future research. Upon publication, the code and datasets will be made publicly accessible.
keywords:
SLAM, Uncrewed surface vehicle, Inland waters, Light detection and ranging (LiDAR), Semantic mapping
[…]
[inst1]organization=Department of Mechanical Engineering,uni=KU Leuven, postcode=3001, city=Leuven, country=Belgium
[…]
[inst2]organization=Department of Computer Science,uni=KU Leuven, postcode=8200, city=Bruges, country=Belgium
[…]
[inst3]organization=Faculty of Mechanical Engineering,uni=TU Eindhoven, postcode=5612 AZ, city=Eindhoven, country=Netherlands
[…]
[inst4]organization=Flanders Make – Leuven,postcode=3001, city=Leuven, country=Belgium
1 Introduction
@@
IWT constitutes an essential component of Europe’s freight infrastructure, spanning a network exceeding 41,000 km, interlinking major cities and industrial hubs across 13 interconnected Member States [1]. As efforts increase to shift freight from congested road and rail networks, the importance of accurate geospatial information and detailed environmental models for managing and navigating these waterways grows [2]. However, navigating these waterways involves safety risks, with studies identifying human error as the primary cause in up to 70% of maritime incidents [3]. In this context, advancements in automation, especially the deployment of uncrewed surface vehicles (USVs), are driving a transformation in inland waterway transport. USVs play a crucial role not only by mitigating the risks associated with human factors, but also by promising improved cargo transport efficiency and facilitating innovative logistics services, including operations within challenging environments such as congested urban canal networks [4]. However, the effective operation of USVs is challenging due to the complex and dynamic nature of inland waterways. Safe and efficient navigation mainly depends on robust systems for perception, localization, and mapping. These systems must enable the vessel to reliably recognize diverse surroundings, including shorelines, infrastructure, vegetation, and overhead obstacles such as bridges, to operate safely.
@@
Traditionally, navigation on inland waterways relies highly on Global Navigation Satellite Systems (GNSS) for positioning, complemented by standard nautical charts like IENC for contextual spatial data [5]. However, GNSS frequently suffer from severe attenuation or complete outages in typical IWT scenarios, particularly during passages under bridges, within narrow urban canals flanked by tall buildings ’urban canyons’, or along waterways with dense vegetation canopies [6, 7]. This documented unreliability highlights the need for onboard localization systems that work independently [8]. Furthermore, existing geospatial data products, IENC, while providing important routing and regulatory data, often lack the spatial detail or real-time updates required for accurate autonomous maneuvering and perception. Representations of shorelines and their surrounding features, including vegetation, shoreline constructions, and other obstacles, are often generalized or outdated in existing maps [9, 10]. A key limitation of IENC is the representation of safety parameters, such as vertical bridge clearance, using static data. This approach fails to account for dynamic water level variations, a significant operational challenge in IWT affecting actual clearances [11]. As a result, relying on static, low-resolution geospatial information hinders the deployment of advanced autonomous functions.
@@
LiDAR-based SLAM has developed as a strong technique in terrestrial robotics and autonomous driving, making it possible to build precise 3D maps and accurate localization even in environments without GNSS [12, 13]. Applying LiDAR SLAM to the IWT is a viable way to overcome the limitations of previous approaches. However, traditional LiDAR SLAM algorithms face particular challenges in this context. The main challenges come from the structure of the environment: unlike urban scenes with abundant planar facades (buildings) and consistent features (lane markings), inland waterways predominantly feature irregular, natural shorelines often covered with vegetation [11]. This lack of structured geometry and standardized visual cues affects the reliable extraction of edges and planar features necessary by algorithms like LOAM [14] and LeGO-LOAM [15]. In addition to these algorithmic challenges, the development and evaluation of SLAM algorithms for IWT is hindered by the lack of suitable benchmark datasets.
@@
Beyond achieving accurate geometric reconstruction, effective environmental understanding demands semantic understanding, e.g. distinguishing navigable water from vegetated banks, identifying a potential docking quay wall, or determining safe passage under a bridge. Standard SLAM methods typically do not provide this layer of structural semantic information. Finally, a gap exists in translating the rich, high-resolution data from SLAM systems into formats aligned with IENC navigation standards.
@@
Figure 1: Our experimental platform, a catamaran named the Maverick, was equipped with a Sensor Box for situational awareness and data collection.
@@
    Inland-LOAM is specifically enhanced for robustness in GNSS-denied inland waterway environments through improved feature extraction performance and the incorporation of planar surface optimization.
@@
    The proposed framework was demonstrated in a diverse set of real-world inland waterway scenes, including ports, urban canals, vegetated banks, bridges, and lifting gates. The experimental platform used for data collection is shown in Fig. 1. All data are released as part of the publicly available dataset to support further research.
2 Related Work
@@
2.1 Mapping in Inland Waterway Environments
Inland waterways are mapped using various methods depending on their scale and purpose. Large-scale remote sensing, using airborne or satellite platforms with LiDAR, multispectral imagery, or synthetic aperture radar (SAR), provides maps and monitoring of waterway networks and adjacent land cover [16, 17]. However, these top-down approaches often yield static data lacking the resolution and perspective needed for capturing critical navigational details. LiDAR-based onboard mapping offers a promising solution for detailed mapping in inland scenarios [18, 19, 20, 21]. However, deploying LiDAR SLAM in IWT presents distinct complexities compared to established terrestrial applications [22, 23]:
@@ 2.1 Mapping in Inland Waterway Environments
2.2 LiDAR Odometry and Mapping and Variants
@@
LOAM [14] pioneered efficient, real-time LiDAR odometry based on the roughness of each point on its respective ring. Its low drift and computational efficiency made it has been ranking top-3 in the KITTI odometry leaderboard [28]. LeGO-LOAM [15] built upon LOAM by incorporating point cloud segmentation to separate ground points from other objects and clustering non-ground points. This allows focusing feature extraction on more salient objects and significantly reduces computational cost, making it suitable for platforms with limited processing power. Subsequent works, such as F-LOAM [29], further focused on developing lightweight LOAM variants with optimized efficiency. For flat-ground scenarios, HDL-Graph-SLAM [30] employs graph optimization using NDT scan matching integrated with ground plane constraints. Ground-SLAM [31] specifically addresses structured multi-floor environments by extracting ground segments and modeling them as planar landmarks via Closest Point (CP) parameterization. Recently, KISS-ICP [32] achieved high performance through simple but effective point-to-point registration and accurate motion compensation. However, methods explicitly incorporating ground plane constraints typically assume the existence of a continuous solid ground. Additionally, ICP-based approaches, despite their computational efficiency, often overlook the richer geometric information provided by feature-based methods.
2.3 Semantic SLAM and Scene Understanding
@@
Semantic SLAM commonly relies on Deep Learning (DL), SuMa++ [33] use Fully Convolutional Networks (FCNs) [34] for segmentation that allows filtering of dynamic objects. Others [35] leverage Transformers [36] for integrated feature extraction and segmentation, enabling selective processing focused on static scene components to enhance localization accuracy. While effective for object recognition, such DL methods typically rely on large labeled datasets.
@@
In contrast, geometric-based approaches derive semantic understanding without extensive annotations. $\mathcal{PLC}$-LiSLAM [37] detects geometric primitives, such as planes, lines, and cylinders, as stable landmarks for SLAM optimization. Similarly, [38] achieves forest localization by clustering to identify trees and matching them to aerial canopy maps. These methods leverage inherent geometric structures for environmental understanding, avoiding the need for learned object appearance models.
2.4 Inland Navigation Standards and Map Integration
@@
IENCs are vital for navigation on European inland waterways, delivering essential details on waterway layouts, aids to navigation, locks, and bridges [39]. Recent studies have explored enhancing IENCs with sensor data. For example, Van Baelen et al. [9] used onboard LiDAR to enhance situational awareness for inland vessels, detecting dynamic environments via a KDTree approach and aligning them with IENC data. Pieper and Hahn [10] proposed a 3D LiDAR sensor fusion method to improve ENC accuracy in harbor areas. Our prior work, Luo [18], dynamically updated IENCs using clustering techniques on LiDAR data. While these efforts highlight the potential of LiDAR for IENC updates, they often depend on direct projection and ad-hoc feature fitting. Similarly, Fredriksson et al. [40] converted voxel maps into 2D occupancy grids for path planning by projecting verified free space and preserving height and slope metadata, suggesting a promising pathway for efficient 3D-to-2D map conversion. However, existing research either validates isolated ENC objects or compresses 3D maps for robotic planning, lacking an end-to-end algorithm to transform LiDAR-derived structural and semantic data into standardized geospatial overlays suitable for IENCs.
3 Methodology
3.1 System Overview
@@
The Inland-LOAM framework, as shown in Fig. 2, is designed to build structural semantic maps of inland waterways for enhanced situational awareness, only relying on LiDAR. The system comprises two primary modules: first, a robust LiDAR Odometry and Mapping (LOAM) pipeline for accurate pose estimation and initial environmental representation; and secondly, a semantic interpretation and conversion module to produce enhanced, IENC-compatible maps.
@@
In the LOAM module, raw LiDAR point clouds are first continuously received by the system. These point clouds then undergo a pre-processing component to remove dynamic points from the vessel itself and to project the data onto a range image, which is a mapping in the angular coordinate, for segmentation. Next, an improved feature extraction method identifies distinct geometric features—such as surface, flat, and rough points. These features are then employed by a two-stage LiDAR Odometry component for pose estimation, which is processed further by a dedicated Plane Fitting component. Finally, the LiDAR mapping component incrementally builds the map and optimizes the trajectory by registering new scans against the global representation, integrating pose graph optimization via incremental smoothing and mapping (iSAM2) [41] within the Georgia Tech Smoothing and Mapping (GTSAM) [42] library to ensure global consistency.
@@
Figure 2: Overview of the Inland-LOAM framework, which processes LiDAR data to generate IENC-compatible structural semantic maps for inland waterways. The LOAM pipeline (upper section) performs robust LiDAR odometry and constructs a globally consistent map. The semantic interpretation and conversion module (lower section) then uses this map to build a Voxel Map, analyze geometric properties for structural semantic understanding, and produce a 2D map converted into an IENC-compatible format.
@@
The output from the LOAM module, comprising the point cloud map and estimated poses, serves as the input for the Conversion module in Fig. 2 lower part. In this module, a Voxel Map is created by dynamically combining features from the LOAM output based on probabilistic occupancy grids [43]. Next, 3D free and occupied space are identified and extracted from this Voxel Map. Slope estimation is then performed based on the extracted occupied regions. These estimated slopes, combined with local neighbourhood features from the Voxel Map, are subsequently used to determine structural semantic attributes, which collectively form a 2D Structural Semantic Map. Finally, a last conversion stage converts this 2D map into lightweight data that can be integrated with standard navigational charts,
3.2 System Equipment
@@
The data acquisition setup, as illustrated in Fig. 3, integrates a Robosense Helios-32 LiDAR with a Septentrio AsteRx-i3 D Pro+ GNSS/INS receiver.
@@
In response to the inherent water-surface interference challenges of inland waterways discussed in Section 2.1 [24, 25], Allis et al. [44] demonstrated that optimal detection of water surfaces relies on diffuse reflectivity, influenced predominantly by laser incidence angles and water turbidity. Furthermore, optical constraints, such as grazing angles, inherently limit the detectable surface area. Therefore, a LiDAR with an expanded vertical field-of-view was selected to maximize the reliability of water surface detection.
@@
3.3 LiDAR Odometry And Mapping Module
3.3.1 Preprocessing and Plane Fitting
The preprocessing begins by converting the raw LiDAR measurements into a structured range image. Let $\mathbf{P}_{k}=\{\mathbf{p}_{i}\}_{i=1}^{N}$ represent the LiDAR point cloud collected in the $k$ th scan cycle, where each point $\mathbf{p}_{i}=[x_{i},y_{i},z_{i}]^{\top}\in\mathbb{R}^{3}$ denotes spatial coordinates. Due to the substantial vertical FoV of the LiDAR and the considerable physical dimensions of the vessel, some of the LiDAR returns inevitably originate from the vessel itself. Therefore, a bounding-box ($\text{bbox}_{\text{vessel}}$) filtering step is applied to exclude these points:

### v2

# Inland-LOAM: Voxel-Based Structural Semantic LiDAR Odometry and Mapping for Inland Waterway Navigation

@@ Abstract
Accurate and up-to-date geospatial information is crucial for enhancing the safety and autonomy of Inland Waterway Transport (IWT). Existing Inland Electronic Navigational Charts (IENCs) often lack real-time detail, and conventional LiDAR-based Simultaneous Localization and Mapping (SLAM) methods struggle with waterway characteristics like unstructured geometric features and water surface reflections. These challenges lead to significant localization drift and produce point cloud maps lacking the semantic richness required for autonomous decision-making. This paper introduces a comprehensive LiDAR odometry and Mapping framework for inland waterway navigation (Inland-LOAM). We present an improved feature extraction method adapted to unique waterway geometries, combined with a joint optimization that incorporates the water surface as a global planar constraint to mitigate drift. We also propose an innovative pipeline that transforms dense 3D point cloud outputs into structured 2D semantic maps. By constructing semantic voxel grids and performing geometric analyses (roughness, planarity, and slope), our system classifies the environment into meaningful structural categories and supports real-time computation of critical parameters like vertical bridge clearances. An automated module then efficiently extracts shoreline boundaries, exporting them into a lightweight, IENC-compatible format. Extensive evaluations on a diverse, real-world dataset demonstrate that Inland-LOAM achieves superior localization accuracy over state-of-the-art methods. The generated maps and shorelines align with real-world conditions, providing reliable information to enhance navigational situational awareness. Both the dataset and the algorithm are publicly available to support future research.
I Introduction
@@
IWT constitutes an essential component of Europe’s freight infrastructure, spanning a network exceeding 41,000 km, interlinking major cities and industrial hubs across 13 interconnected Member States [1]. As efforts increase to shift freight from congested road and rail networks, the importance of accurate geospatial information and detailed environmental models for managing and navigating these waterways grows [2]. However, navigating these waterways involves safety risks, with studies identifying human error as the primary cause in up to 70% of maritime incidents [3]. In this context, advancements in automation, especially the deployment of uncrewed surface vehicles (USVs), are driving a transformation in inland waterway transport. USVs play a crucial role not only by mitigating the risks associated with human factors, but also by promising improved cargo transport efficiency and facilitating innovative logistics services, including operations within challenging environments such as congested urban canal networks [4]. However, the effective operation of USVs is challenging due to the complex and dynamic nature of inland waterways. These are confined channels with locks and (movable and fixed) bridges and limited draft, which often create navigation bottlenecks [5]. Safe and efficient navigation mainly depends on robust systems for perception, localization, and mapping. These systems must enable the vessel to reliably recognize diverse surroundings, including shorelines, infrastructure, vegetation, and overhead obstacles such as bridges, to operate safely.
@@
Traditionally, navigation on inland waterways relies highly on Global Navigation Satellite Systems (GNSS) for positioning, complemented by standard nautical charts like IENC for contextual spatial data [6]. However, GNSS frequently suffers from severe attenuation or complete outages in typical IWT scenarios, particularly during passages under bridges, within narrow urban canals flanked by tall buildings ”urban canyons”, or along waterways with dense vegetation canopies [7, 8]. This documented unreliability highlights the need for onboard localization systems that work independently [9]. Furthermore, existing geospatial data products, IENC, while providing important routing and regulatory data, often lack the spatial detail or real-time updates required for accurate autonomous maneuvering and perception. Representations of shorelines and their surrounding features, including vegetation, shoreline constructions, and other obstacles, are often generalized or outdated in existing maps [10, 11]. A key limitation of IENC is the representation of safety parameters, such as vertical bridge clearance, using static data. This approach fails to account for dynamic water level variations, a significant operational challenge in IWT affecting actual clearances [12]. As a result, relying on static, low-resolution geospatial information hinders the deployment of advanced autonomous functions.
@@
LiDAR-based SLAM has developed as a strong technique in terrestrial robotics and autonomous driving, making it possible to build precise 3D maps and accurate localization even in environments without GNSS [13, 14]. Applying LiDAR SLAM to the IWT is a viable way to overcome the limitations of previous approaches. However, traditional LiDAR SLAM algorithms face particular challenges in this context. The main challenges come from the structure of the environment: unlike urban scenes with abundant planar facades (buildings) and consistent features (lane markings), inland waterways predominantly feature irregular, natural shorelines often covered with vegetation [12]. This lack of structured geometry and standardized visual cues affects the reliable extraction of edges and planar features necessary by algorithms like LOAM [15] and LeGO-LOAM [16]. Although this lack of structured geometry presents a challenge, these environments simultaneously offer a unique and powerful geometric prior that is often absent in terrestrial scenarios: the water surface itself. Rather than being treated as a source of interference, we contend that when properly detected, the inland water surface provides a vast, naturally occurring, and globally referenced planar constraint. This offers a simple, yet robust anchor to mitigate drift, particularly for roll, pitch, and altitude - a benefit not available from the smaller, localized planes found in typical indoor or urban environments. In addition to these algorithmic challenges, the development and evaluation of SLAM algorithms for IWT is hindered by the lack of suitable benchmark datasets.
@@
Beyond achieving accurate geometric reconstruction, effective environmental understanding demands semantic understanding, e.g., distinguishing navigable water from vegetated banks, identifying a potential docking quay wall, or determining safe passage under a bridge. Standard SLAM methods typically do not provide this layer of structural semantic information. Finally, a gap exists in translating the rich, high-resolution data from SLAM systems into formats aligned with IENC navigation standards.
@@
Figure 1: Our experimental platform, a catamaran named the Maverick, was equipped with a Sensor Box for situational awareness and data collection. Image source: [17].
@@
    Inland-LOAM is specifically enhanced for robustness in GNSS-denied inland waterway environments through improved feature extraction performance and the incorporation of water surface optimization to mitigate the prevalent vertical drift in inland waterway navigation.
@@
    The proposed framework was demonstrated in a diverse set of real-world inland waterway scenes, including ports, urban canals, vegetated banks, bridges, and lifting gates. The experimental platform used for data collection is shown in Figure 1. All data are released as part of the publicly available dataset to support further research.
II Related Work
@@
II-A Mapping in Inland Waterway Environments
LiDAR-based onboard mapping offers a promising solution for detailed mapping in inland scenarios [18, 19, 20, 21]. However, deploying LiDAR SLAM in IWT presents distinct complexities compared to established terrestrial applications [22, 23]:
@@ II-A Mapping in Inland Waterway Environments
II-B LiDAR Odometry and Mapping and Variants
@@
LOAM [15] pioneered efficient, real-time LiDAR odometry based on the roughness of each point on its respective ring. Its low drift and computational efficiency placed it among the top-3 methods on the KITTI odometry benchmark [28]. LeGO-LOAM [16] built upon LOAM by incorporating point cloud segmentation to separate ground points from other objects and clustering non-ground points. This allows focusing feature extraction on more salient objects and significantly reduces computational cost, making it suitable for platforms with limited processing power. Subsequent works, such as F-LOAM [29], further focused on developing lightweight LOAM variants with optimized efficiency. For flat-ground scenarios, HDL-Graph-SLAM [30] employed graph optimization using NDT scan matching integrated with ground plane constraints. Ground-SLAM [31] specifically addressed structured multi-floor environments by extracting ground segments and modeling them as planar landmarks via Closest Point (CP) parameterization. Recently, KISS-ICP [32] achieved high performance through simple but effective point-to-point registration and accurate motion compensation. However, methods explicitly incorporating ground plane constraints typically assume the existence of a continuous solid ground. Additionally, ICP-based approaches, despite their computational efficiency, often overlook the richer geometric information provided by feature-based methods.
II-C Semantic SLAM and Scene Understanding
@@
Semantic SLAM commonly relies on Deep Learning (DL). SuMa++ [33] used Fully Convolutional Networks (FCNs) [34] for segmentation that allows filtering of dynamic objects. Similarly, Gong et al. [35] integrated LiDAR SLAM with a PointNet-based segmentation head to obtain semantic modeling in GNSS-denied underground parking lots. Other [36] leveraged Transformers [37] for integrated feature extraction and segmentation, enabling selective processing focused on static scene components to enhance localization accuracy. While effective for object recognition, such DL methods typically rely on large labeled datasets.
@@
In contrast, geometric-based approaches derive semantic understanding without extensive annotations. $\mathcal{PLC}$-LiSLAM [38] detected geometric primitives, such as planes, lines, and cylinders, as stable landmarks for SLAM optimization. Similarly, [39] achieved forest localization by clustering to identify trees and matching them to aerial canopy maps. These methods leverage inherent geometric structures for environmental understanding, avoiding the need for learned object appearance models.
II-D Inland Navigation Standards and Map Integration
@@
IENCs are vital for navigation on European inland waterways, delivering essential details on waterway layouts, aids to navigation, locks, and bridges [40]. Recent studies have explored enhancing IENCs with sensor data. For example, Van Baelen et al. [10] used onboard LiDAR to enhance situational awareness for inland vessels, detecting dynamic environments via a KDTree approach and aligning them with IENC data. Pieper and Hahn [11] proposed a 3D LiDAR sensor fusion method to improve ENC accuracy in harbor areas. Our prior work, Luo [18], dynamically updated IENCs using clustering techniques on LiDAR data. While these efforts highlight the potential of LiDAR for IENC updates, they often depend on direct projection and ad-hoc feature fitting. Similarly, Fredriksson et al. [41] converted voxel maps into 2D occupancy grids for path planning by projecting verified free space and preserving height and slope metadata, suggesting a promising pathway for efficient 3D-to-2D map conversion. However, existing research either validates isolated ENC objects or compresses 2D maps for robotic planning, lacking an end-to-end algorithm to transform LiDAR-derived structural and semantic data into standardized geospatial overlays suitable for IENCs.
III Methodology
III-A System Overview
@@
The Inland-LOAM framework, as shown in Figure 2, is designed to build structural semantic maps of inland waterways for enhanced situational awareness, only relying on LiDAR. The system comprises two primary modules: first, a robust LiDAR Odometry and Mapping (LOAM) pipeline for accurate pose estimation and initial environmental representation; and secondly, a semantic interpretation and conversion module to produce enhanced, IENC-compatible maps.
@@
In the LOAM module, raw LiDAR point clouds are first continuously received by the system. These point clouds then undergo a pre-processing component to remove dynamic points from the vessel itself and to project the data onto a range image, which is a mapping in the angular coordinate, for segmentation. Next, an improved feature extraction method identifies distinct geometric features—such as surface, flat, and rough points. These features are then employed by a two-stage LiDAR Odometry component for pose estimation, which is processed further by a dedicated Plane Fitting component. Finally, the LiDAR mapping component incrementally builds the map and optimizes the trajectory. This is formulated as a factor graph optimization problem, which is solved incrementally using iSAM2 (incremental smoothing and mapping) [42] within GTSAM (the Georgia Tech Smoothing and Mapping) [43] library.
@@
Figure 2: Overview of the Inland-LOAM framework, which processes LiDAR data to generate IENC-compatible structural semantic maps for inland waterways. The LOAM pipeline (upper section) performs robust and low-drift LiDAR odometry. The semantic interpretation and conversion module (lower section) then uses this map to build a Voxel Map, analyze geometric properties for structural semantic understanding, and produce a 2D map converted into an IENC-compatible format.
@@
The output from the LOAM module, comprising the point cloud map and estimated poses, serves as the input for the Conversion module in Figure 2 lower part. In this module, a Voxel Map is created by dynamically combining features from the LOAM output based on probabilistic occupancy grids [44]. Next, 3D free and occupied space are identified and extracted from this Voxel Map. Slope estimation is then performed based on the extracted occupied regions. These estimated slopes, combined with local neighbourhood features from the Voxel Map, are subsequently used to determine structural semantic attributes, which collectively form a 2D Structural Semantic Map. Finally, a last conversion stage converts this 2D map into lightweight data that can be integrated with standard navigational charts.
III-B System Equipment
@@
The data acquisition setup, as illustrated in Figure 3, integrates a Robosense Helios-32 LiDAR with a Septentrio AsteRx-i3 D Pro+ GNSS/INS receiver.
@@
In response to the inherent water-surface interference challenges of inland waterways discussed in Section II-A, Allis et al. [45] demonstrated that optimal detection of water surfaces relies on diffuse reflectivity, influenced predominantly by laser incidence angles and water turbidity. Furthermore, optical constraints, such as grazing angles, inherently limit the detectable surface area. Therefore, a LiDAR with an expanded vertical field-of-view was selected to maximize the reliability of water surface detection.
@@
III-C LiDAR Odometry And Mapping Module
III-C1 Preprocessing and Plane Fitting
The preprocessing begins by converting the raw LiDAR measurements into a structured range image. Let $\mathbf{P}_{k}=\{\mathbf{p}_{i}\}_{i=1}^{N}$ represent the LiDAR point cloud collected in the $k$ th scan cycle, where each point $\mathbf{p}_{i}=[x_{i},y_{i},z_{i}]^{\mathsf{T}}\in\mathbb{R}^{3}$ denotes spatial coordinates. Due to the substantial vertical FoV of the LiDAR and the considerable physical dimensions of the vessel, some of the LiDAR returns inevitably originate from the vessel itself. Therefore, a bounding-box ($\text{bbox}_{\text{vessel}}$) filtering step is applied to exclude these points: